# Retrieving snow depth distribution by downscaling ERA5 Reanalysis with ICESat-2 laser altimetry

### v1

@@ Abstract
Estimating the variability of seasonal snow cover, in particular snow depth in remote areas, poses significant challenges due to limited spatial and temporal data availability. This study uses snow depth measurements from the ICESat-2 satellite laser altimeter, which are sparse in both space and time, and incorporates them with climate reanalysis data into a downscaling-calibration scheme to produce monthly gridded snow depth maps at microscale (10 m). Snow surface elevation measurements from ICESat-2 along profiles are compared to a digital elevation model to determine snow depth at each point. To efficiently turn sparse measurements into snow depth maps, a regression model is fitted to establish a relationship between the retrieved snow depth and the corresponding ERA5 Land snow depth. This relationship, referred to as subgrid variability, is then applied to downscale the monthly ERA5 Land snow depth data. The method can provide timeseries of monthly snow depth maps for the entire ERA5 time range (since 1950). We observe that the generic output should be calibrated by a small number of localized control points from a one-time field survey to restore the full snow depth patterns. The validation of downscaled snow depth data was performed at an intermediate scale (100 m x 500 m) using datasets from airborne laser scanning (ALS) in the Hardangervidda region of southern Norway. Results show that snow depth prediction achieved $R^{2}$ values ranging from 0.74 to 0.88 (post-calibration). The method relies on globally available data and is applicable to other snow regions above the treeline. Though requiring area-specific calibration, our approach has the potential to provide snow depth maps in areas where no such data exist and can be used to extrapolate existing snow surveys in time and over larger areas. With this, it can offer valuable input data for hydrological, ecological or permafrost modeling tasks.
@@
[…]
[1]organization=University of Oslo, Department of Geosciences,addressline=Sem Sælands vei 1,city=Oslo,postcode=371,postcodesep=
@@ 1 Introduction
In a warming world, understanding the spatio-temporal variations of seasonal snow is increasingly vital for climate impact assessments, meltwater supply [36, 47], permafrost modeling [23] and ecological responses [6]. Seasonal snow accumulates and melts away once a year in a number of climatic zones from, for instance, the forested regions of the taiga (largest terrestrial ecosystem), the open tundra of the Arctic and many high mountain ranges throughout the world [79]. Being white, snow is easily observable from space. However, its thin nature (typically less than 2 m) makes it very challenging to measure snow thickness or mass from space at a large scale and with high repeatability. The complex processes driving snow metamorphism and precipitation estimation are hurdles too to rely on models for assessing seasonal snow dynamics. Observing and modeling seasonal snow at a large scale therefore remains a challenge [86, 63]. This is exacerbated in remote and complex terrain with limited data availability [4].
@@
A major persistent gap is observing or estimating snow depth and water equivalent over mountain regions. In-situ point-based snow stations are generally located in gentle terrain at lower or mid-elevation, often under-representing rugged and higher elevation [18]. Employing sensors on airplanes or UAVs (unmanned aircraft vehicles) offers very high-resolution data [5], but are costly, of limited spatial extent and potentially with legal flight restrictions. However this can be an effective approach for watershed-scale snow depth mapping [12].
@@
Otherwise, space-borne technologies can cover systematically large to global extents at regular intervals. Two main technologies exist to retrieve snow depth: radar and laser altimetry. Radar has several issues with footprint size, penetration depth, snow wetness, and modeling effects due to scattering [14]. For instance, passive microwave radar with a footprint size of up to 25 km is insufficient to capture the heterogeneous snow depth in mountainous terrain [86, 63]. Despite recent technology development in C-band radar enabling snow depth retrieval at resolutions ranging from 500 meters to 1 kilometer under dry snow conditions [44, 43], finer-scale snow depth remains poorly addressed (e.g. 100 m resolution, Grünewald et al. 24, Mott et al. 60). Another approach for high-resolution snow depth data is to use laser scanning or photogrammetry. Both methods provide snow depth by collecting surface height data pre- and post-snowfall, enabling snow depth mapping through comparative analysis of height differences. Recently, there has been a significant effort to retrieve snow depth by combining snow-free digital elevation models (DEMs) with space-borne lidar measurements of the snow surface from ICESat [83] and its successor ICESat-2 [66]. This cutting-edge satellite offers high-resolution, accurate lidar elevation profile measurements of the Earth’s surface, including snow-covered terrains.
@@
Deschamps-Berger et al. [13] derived snow depth from the ICESat-2 ATL06 products and reported an accuracy of 0.2 m (bias) and a precision (normalized median absolute deviation; NMAD) of 0.5 m for low slopes and 1.2 m for steeper areas over the upper Tuolumne basin, California, USA. Enderlin et al. [16] found that snow depth estimates based on ICESat-2 data had a median absolute deviation (MAD) ranging from 0.2 m for slopes < 5° to over 1 m for slopes > 20°. Besso et al. [3] questioned the varying accuracy of the ICESat-2 ATL08 product and developed a self-defined processed elevation product, which achieved a MAD of 0.14 m to 0.20 m and root mean square error (RMSE) of 0.18 m to 0.33 m for the Tuolumne Basin and Methow Valley, USA. These studies have uncovered ICESat-2 as an emerging and cost-efficient data source for snow depth and also brought attention to challenges associated with the data and elevation differencing workflow, primarily stemming from discrepancies and spatially/temporally varying inconsistencies between reference DEMs and ICESat-2. Therefore, implementing and improving this workflow requires careful co-registration and bias correction on DEMs & ICESat-2. Additionally, the sparse nature of ICESat-2 measurements presents another significant challenge to creating comprehensive snow-depth maps: how can we extrapolate both spatially and temporally to areas outside of the ICESat-2 measurement profiles?
@@
Another approach to understanding snow dynamics is through snow modeling. Researchers primarily use two modeling strategies to study these dynamics, namely process-based and statistical approaches. Process-based models [40, 46, 39] incorporate physical processes, which are driven by meteorological forcing data and yield gridded snow depth products. Mazzolini et al. [52] combine snow depth transects from the high-resolution ICESat-2 ATL03 product with snow modelling in a data assimilation framework. They spatially propagate sparse ICESat-2 snow profile information using an abstract distance measured in a feature space defined by topographical parameters and the snow melt-out climatology. For the 1 km2 Izas catchment in the Spanish Pyrenees, they show that adding snow depth information in addition to the traditionally used fractional snow-covered area observations improves the model skill score by 22%. However, these models are computationally costly and thus often struggle to cover large areas or offer fine resolution, and are hindered by complex near-surface atmospheric processes and limited data on precipitation and wind fields [21], leading to a new question: how to quantify the subgrid variability of snow depth [9]. The distribution patterns of snow exhibit a notable resemblance year after year due to their dependence on topography, vegetation, and consistent synoptic weather patterns [80, 71]. The consistent recurrence of this pattern supports the use of computationally efficient statistical approaches. Many studies seek to establish parameterizations for subgrid variability, such as snow depletion curves, snow depth elevation gradients, snow probability distribution[57], subgrid snow depth coefficient of variation [45, 29, 23] or topographic correlations [31, 52]. Learning and reproducing a snow depth map at fine scales typically involves pattern recognition. Multiple-linear regression [25, 15], binary regression trees [75], random forests [74] or a convolutional neural network [11] have been used to predict snow distribution patterns with varied performance ($R^{2}$ of 0.25-0.91). However, these statistical models typically require substantial training data from terrestrial or airborne sensors. Therefore, most models can hardly be transferred to other catchments or seasons [25, 74]. Another category of statistical models capable of generalizing subgrid variability is commonly known as downscaling models. These models are designed to refine data from coarse, broad-scale grids to localized subgrid levels [50]. When applying these models retrospectively or into the future, an important assumption is made: that the statistical relationships remain constant over time, a condition known as stationarity. Currently, there is a limited number of statistical downscaling models applied to snow depth [30, 85]. The primary obstacles involve obtaining sufficient snow depth measurements for training and testing, meteorological forcing data in high resolution and accurately recognising variability through informative features.
@@
The study area of Norway/Hardangervidda mountain plateau is chosen based on the availability of validation data. Still, the workflow is designed to be globally applicable where an accurate DEM and a proper calibration dataset are available. To retrieve snow depth measurements from ICESat-2 data, we use nationally and globally available DEMs acquired during snow-free conditions. Subsequently, we statistically downscaled ERA5 Land from its ~9 km native resolution into 10 m using ICESat-2 snow depth measurements and a machine learning algorithm. To our knowledge, this study marks the first attempt to use ICESat-2 data to downscale ERA5 Land data, and the first attempt to propagate ICESat-2 snow depths in time and space using a statistical approach..
@@ 2 Study Area and Data Setting
Figure 1: Map of the study area. Left panel: Hardangervidda mountain plateau showing ICESat-2 tracks, validation data from the 2008/2009 snow survey and available weather stations. ICESat-2 ATL08 snow-on data from October 2018 to October 2022 are displayed in red. The blue ticks at the bottom highlight the ground tracks acquired during April (peak snow). The right panel shows location of the validation area and the total number of available snow-on data segments for mainland Norway.
Norway, located on the western side of the Scandinavian Peninsula in northwestern Europe, spans latitudes from 57° 58’N to 71° 11’ N and longitudes from 4° 40’ E to 30° 58’ E. The country features a diverse topography, ranging from lowland valleys to high mountains. In winter, westerly winds bring moisture, resulting in substantial snowfall from the coast to inland areas. This snow acts as a crucial reservoir for hydropower, emphasizing the importance of estimating snow mass in mountainous environments.
@@ 2.1 ICESat-2 ATL08 elevation data
Launched in September 2018, ICESat-2 is equipped with the Advanced Topographic Laser Altimeter System (ATLAS), which provides photon-counting lidar measurements at a global scale [66]. ATLAS comprises three parallel beam pairs separated by 3.3 km on the ground. Each beam pair includes a strong and weak beam separated by 90 m. The ATLAS system emits a pulse every 0.7 m along the track, covering a circular footprint with a diameter of ~15 m. At mid-latitudes, ICESat-2 ground tracks are generally not repeated but shifted for each overpass to maximise global spatial coverage. The ICESat-2 ATL08 product [level L3A, version 5, 68] offers elevation data in fixed segment sizes of 100 meters along the ground track [67]. For each segment, there are 5 geolocations (subsegments) in 20 m intervals. Instead of using the mean elevation of the 100 m segment [h_te_mean, e.g., 16], we utilized the subsegment height referring to best-fit terrain elevation at the midpoint location of the segment (h_te_best_fit_20m_2). The sub-segment heights are determined through polynomial fitting to terrain photons with slope correction and weighting [67]. Norway has a total of 3,968 ATL08 data granules available for analysis from 14 October 2018, to 12 October 2022. After removing invalid data, and excluding permanent ice and inland water, our dataset consists of 13,197,376 segments, including 4,778,904 snow-free segments on land and 8,418,472 segments with snow cover over land. The coverage of snow-on segments is displayed in Figure 1 with blue ticks representing the ground tracks from April of one of the four years (as ground tracks are not repeated).
@@ 2.2 Snow-off elevation data
As reference ground for snow depth retrieval, we employ the Norwegian DTM1 elevation model [DTM1, 37], a 1 m lidar-based product acquired by Kartverket between 2016 and 2022. As a sensitivity test of DEM resolution, we also utilize the 10-meter resolution variant from the same data provider, DTM10 [37]. To demonstrate the workflow’s applicability in areas without lidar-based elevation products, we incorporate global DEMs such as Copernicus GLO-30 [17], hereafter referred to as COP30 and FABDEM [28, Forest And Buildings Removed Copernicus DEM, hereafter referred to as FAB] as reference ground. COP30 is a 30-meter-resolution Digital Surface Model (DSM) acquired between December 2010 and January 2015 through synthetic aperture radar interferometry (TanDEM-X mission). FAB, a variant of COP30, eliminates buildings and trees using the random forest algorithm, enhancing accuracy. FAB serves as a reference for comparison with COP30.
@@ 2.3 Large-scale reanalysis data
ERA5 Land hourly data (version 5) [64] is an ECMWF (European Centre for Medium-Range Weather Forecasts) reanalysis product covering the period from 1950 to the present. It describes water and energy cycles over global land areas with over 50 variables at a spatial resolution of approximately 9 km [62]. This reanalysis data supplies the necessary forcing data for the downscaling model to generate sub-grid products while also accounting for input errors in the model [27, 72]. ERA5 Land’s snow depth data (sde) represents an instantaneous average of the snow thickness on the ground, excluding snow on vegetation canopy [64]. Additionally, the ERA5 Land monthly [65] dataset contributes instantaneous wind fields (u10, v10) at 10 m above the land surface.

### v2

@@ Abstract
Estimating the variability of snow depth in remote areas poses significant challenges due to limited spatial and temporal data availability. This study uses snow depth measurements from the ICESat-2 satellite laser altimeter, which are sparse in both space and time, and incorporates them with climate reanalysis data into a downscaling-calibration scheme to produce monthly gridded snow depth maps at microscale (10 m). Snow surface elevation measurements from ICESat-2 along profiles are compared to a digital elevation model to determine snow depth at each point. To efficiently turn sparse measurements into snow depth maps, a regression model is fitted to establish a relationship between the retrieved snow depth and the corresponding ERA5 Land snow depth. This relationship, referred to as subgrid variability, is then applied to downscale the monthly ERA5 Land snow depth data. The method can provide timeseries of monthly snow depth maps for the entire ERA5 time range (since 1950). We observe that the generic output should be calibrated by a small number of localized control points from a one-time field survey to reproduce the full snow depth patterns. Results show that snow depth prediction achieved a $R2$ model fit value of 0.81 (post-calibration) at an intermediate scale (100 m x 500 m) using datasets from airborne laser scanning (ALS) in the Hardangervidda region of southern Norway, with still good results at microscale ($R2$ 0.34, RMSE 1.28m, post-calibration). Bias is greatest for extremes, with very deep/shallow snow depths being under- and overestimated, respectively. Snow depth time series modeled at the site level have a slightly smaller RMSE than ERA5 Land data, but are still consistently biased compared to measurements from meteorological stations. Despite such localised bias and a tendency towards average snow depths the model reproduces the relative snow distribution pattern very accurately, both for peak snow (Spearman’s $\rho$ 0.77) and patchy snow meltout in late spring (Matthews correlation coefficient 0.35). The method relies on globally available data and is applicable to other snow regions above the treeline. Though requiring area-specific calibration, our approach has the potential to provide snow depth maps in areas where no such data exist and can be used to extrapolate existing snow surveys in time and over larger areas. With this, it can offer valuable input data for hydrological, ecological or permafrost modeling tasks.
@@ 1 Introduction
In a warming world, understanding the spatio-temporal variations of seasonal snow is increasingly important for climate impact assessments, meltwater supply [39, 52], permafrost modeling [26] and ecological responses [7]. Seasonal snow accumulates and melts away once a year in a number of climatic zones from, for instance, the forested regions of the taiga (largest terrestrial ecosystem), the open tundra of the Arctic and many high mountain ranges throughout the world such as the Alps, the Himalayas the Andies [89]. Being white, snow is easily observable from space. However, its thin nature (typically less than 2 m) makes it challenging to measure snow thickness or mass from space at a large scale and with high repeatability. The complex processes driving snow metamorphism and precipitation estimation are hurdles too to rely on models for assessing seasonal snow dynamics. Observing and modeling seasonal snow at a large scale therefore remains a challenge [97, 69]. This is exacerbated in remote and complex terrain with limited data availability [5].
@@
A major persistent gap is observing or estimating snow depth and water equivalent over mountain regions. In-situ point-based snow stations are generally located in gentle terrain at lower or mid-elevation, often under-representing rugged and higher elevation [20]. Employing sensors on airplanes or UAVs (unmanned aircraft vehicles) offers very high-resolution data [6, 77, 79], but are costly, of limited spatial extent and potentially limited by legal flight restrictions. However this can be an effective approach for watershed-scale snow depth mapping [14].
@@
Otherwise, space-borne technologies can cover systematically large to global extents at regular intervals. Two main technologies exist to retrieve snow depth: radar and laser altimetry. Radar has several issues with footprint size, penetration depth, snow wetness, and modeling effects due to scattering [16]. For instance, passive microwave radar with a footprint size of up to 25 km is insufficient to capture the heterogeneous snow depth in mountain terrain [97, 69]. Despite recent technology development in C-band radar enabling snow depth retrieval at resolutions ranging from 500 meters to 1 kilometer under dry snow conditions [48, 47], finer-scale snow depth retrieval remains poorly addressed (e.g. 100 m resolution, Grünewald et al. 27, Mott et al. 66). Another approach for high-resolution snow depth retrieval is to use laser scanning or photogrammetry. Both methods provide snow depth by collecting surface height data pre- and post-snowfall, enabling snow depth mapping by differencing elevations. Recently, there has been a significant effort to retrieve snow depth by combining snow-free digital elevation models (DEMs) with space-borne lidar measurements of the snow surface from ICESat [94] and its successor ICESat-2 [72]. This cutting-edge satellite offers high-resolution, accurate lidar elevation profile measurements of the Earth’s surface, including snow-covered terrains.
@@
Deschamps-Berger et al. [15] derived snow depth from the ICESat-2 ATL06 products and reported an accuracy of 0.2 m (bias) and a precision (normalized median absolute deviation; NMAD) of 0.5 m for low slopes and 1.2 m for steeper areas over the upper Tuolumne basin, California, USA. Enderlin et al. [18] found that snow depth estimates based on ICESat-2 data had a median absolute deviation (MAD) ranging from 0.2 m for slopes < 5° to over 1 m for slopes > 20°. Besso et al. [4] questioned the varying accuracy of the ICESat-2 ATL08 product and developed a self-defined processed elevation product, which achieved a MAD of 0.14 m to 0.20 m and root mean square error (RMSE) of 0.18 m to 0.33 m for the Tuolumne Basin and Methow Valley, USA. These studies have uncovered ICESat-2 as an emerging and cost-efficient data source for snow depth and also brought attention to challenges associated with the data and elevation differencing workflow, primarily stemming from discrepancies and spatially/temporally varying inconsistencies between reference DEMs and ICESat-2. Therefore, implementing and improving this workflow requires careful co-registration and bias correction on DEMs & ICESat-2. Additionally, the sparse nature of ICESat-2 measurements presents another significant challenge to deriving comprehensive snow-depth maps: how can we extrapolate both spatially and temporally to areas outside of the ICESat-2 measurement profiles?
@@
Another approach to understanding snow dynamics is through snow modeling. Researchers primarily use two modeling strategies to study these dynamics, namely process-based and statistical approaches. Process-based models [44, 50, 98, 42] incorporated physical processes, which are driven by meteorological forcing data and yield gridded snow depth products. Mazzolini et al. [57] combined snow depth transects from the high-resolution ICESat-2 ATL03 product with snow modelling in a data assimilation framework. They spatially propagated sparse ICESat-2 snow profile information using an abstract distance measured in a feature space defined by topographical parameters and snow melt-out climatology. For the 1 km2 Izas catchment in the Spanish Pyrenees, they show that adding snow depth information in addition to the traditionally used fractional snow-covered area observations improves the model skill score by 22%. However, these models are computationally costly and thus often struggle to cover large areas or provide fine resolution. Models are hindered by complex near-surface atmospheric processes and limited data on precipitation and wind fields [23], leading to a new question: how to quantify the subgrid variability of snow depth [10]? The distribution patterns of snow exhibit a notable repeatability year after year due to their dependence on topography, vegetation, and consistent synoptic weather patterns [90, 80]. The consistent recurrence of this pattern supports the use of computationally efficient statistical approaches. Many studies seek to establish parameterizations for subgrid variability, such as snow depletion curves, snow depth elevation gradients, snow probability distribution[62], subgrid snow depth coefficient of variation [49, 32, 26] or topographic correlations [34, 57]. Learning and reproducing a snow depth map at fine scales typically involves pattern recognition. Multiple-linear regression [28, 17], binary regression trees [84], random forests [83] or a convolutional neural network [13] have been used to predict snow distribution patterns with varied performance ($R2$ of 0.25-0.91). However, these statistical models typically require substantial training data from terrestrial or airborne sensors. Therefore, most models can hardly be transferred to other catchments or seasons [28, 83]. Another category of statistical models capable of generalizing subgrid variability is commonly known as downscaling models. These models are designed to refine data from coarse, broad-scale grids to localized subgrid levels [55]. When applying these models retrospectively or into the future, an important assumption is made: the statistical relationships remain constant over time, a condition known as stationarity. Currently, there is a limited number of statistical downscaling models applied to snow depth [33, 96]. The primary obstacles involve obtaining sufficient snow depth measurements for training and testing, meteorological forcing data in high resolution and accurately recognising variability through informative features.
@@
The study area of Norway/Hardangervidda mountain plateau is chosen based on the availability of validation data. Still, the workflow is designed to be globally applicable where an accurate DEM and a proper calibration dataset are available. To retrieve snow depth measurements from ICESat-2 data, we used nationally and globally available DEMs acquired during snow-free conditions. Subsequently, we statistically downscaled ERA5 Land from its ~9 km native resolution into 10 m using ICESat-2 snow depth measurements and a machine learning algorithm. To our knowledge, this study marks the first attempt to use ICESat-2 data to downscale ERA5 Land data, and the first attempt to propagate ICESat-2 snow depths in time and space using a statistical approach.
@@ 2 Study Area and Data Setting
Figure 1: Map of the study area, Hardangervidda mountain plateau. A photo of the typical summer landscape is displayed at the top (a), looking south towards the Finse area and Hardangervidda. Photographer: Simon Filhol). The main panel (b) shows ICESat-2 tracks, validation data from the 2008/2009 snow survey and available weather stations. ICESat-2 ATL08 snow-on data from October 2018 to October 2022 are displayed in red. The majority of tracks was surveyed only once and the few repeated tracks appear darker due to higher density. The blue ticks at the bottom highlight the ground tracks acquired during April (peak snow). The right panel shows the location of the validation area and the total number of available snow-on data segments for mainland Norway (c).
Norway, located on the western side of the Scandinavian Peninsula in northwestern Europe, spans latitudes from 57° 58’N to 71° 11’ N and longitudes from 4° 40’ E to 30° 58’ E. The country features a diverse topography, ranging from lowland valleys to high mountains (highest peak at 2469 m a.s.l.). In winter, westerly winds bring moisture, resulting in substantial snowfall from the coast to inland areas. This snow acts as a crucial reservoir for hydropower, emphasizing the importance of estimating snow mass in mountain environments.
@@ 2.1 ICESat-2 ATL08 elevation data
Launched in September 2018, ICESat-2 is equipped with the Advanced Topographic Laser Altimeter System (ATLAS), which provides photon-counting lidar measurements at a global scale [72]. ATLAS comprises three parallel beam pairs separated by 3.3 km on the ground. Each beam pair includes a strong and weak beam separated by 90 m. The ATLAS system emits a pulse every 0.7 m along the track, covering a circular footprint with a diameter of ~15 m. At mid-latitudes, ICESat-2 ground tracks are generally not repeated but shifted for each overpass to maximise global spatial coverage. The ICESat-2 ATL08 product [level L3A, version 5, 74] offers elevation data in fixed segment sizes of 100 meters along the ground track [73]. For each segment, there are five geolocations (subsegments) in 20 m intervals. Instead of using the mean elevation of the 100 m segment [h_te_mean, e.g., 18], we used the subsegment height referring to best-fit terrain elevation at the midpoint location of the segment (h_te_best_fit_20m_2). The sub-segment heights are determined through polynomial fitting to terrain photons with slope correction and weighting [73]. Norway has a total of 3,968 ATL08 data granules available for analysis from 14 October 2018, to 12 October 2022. After removing invalid data, and excluding permanent ice and inland water, our dataset consists of 13,197,376 segments, including 4,778,904 snow-free segments on land and 8,418,472 segments with snow cover over land. The coverage of snow-on segments is displayed in Figure 1 with blue ticks representing the ground tracks from April of one of the four years (as ground tracks are not repeated).
@@ 2.2 Snow-off elevation data
As reference ground for snow depth retrieval, we employ the Norwegian DTM1 elevation model [DTM1, 40], a 1 m lidar-based product acquired by Kartverket between 2016 and 2022. As a sensitivity test of DEM resolution, we also use the 10-meter resolution variant from the same data provider, DTM10 [40]. To demonstrate the workflow’s applicability in areas without lidar-based elevation products, we incorporate global DEMs such as Copernicus GLO-30 [19], hereafter referred to as COP30, and FABDEM [31, Forest And Buildings Removed Copernicus DEM, hereafter referred to as FAB] as reference ground. COP30 is a 30-meter-resolution Digital Surface Model (DSM) acquired between December 2010 and January 2015 through synthetic aperture radar interferometry (TanDEM-X mission). FAB, a variant of COP30, eliminates buildings and trees using the random forest algorithm, enhancing accuracy. FAB serves as a reference for comparison with COP30.
@@ 2.3 Large-scale reanalysis data
ERA5 Land hourly data (version 5) [70] is an ECMWF (European Centre for Medium-Range Weather Forecasts) reanalysis product covering the period from 1950 to the present. It describes water and energy cycles over global land areas with over 50 variables at a spatial resolution of approximately 9 km [68]. This reanalysis data supplies the necessary forcing data for the downscaling model to generate sub-grid products while also accounting for input errors in the model [30, 81]. ERA5 Land’s snow depth data (sde) represents the instantaneous snow thickness on the ground for the elevation of each grid cell, excluding snow on vegetation canopy [70]. Additionally, the ERA5 Land monthly [71] dataset contributes instantaneous wind fields (u10, v10) at 10 m above the land surface.